# El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data

### v1

@@ 3 Stellar kinematics measurements
Figure 2: Extraction of MUSE aperture spectra within circular apertures, for L1, L2, L3 and the western part of El Anzuelo arc. Top left panel: luminosity-weighted median stack of the MUSE datacube, with contours from NIRCam/F [^4] W observations. Bottom left panel: MUSE median stack with circular apartures used to extract spectra. Circular apertures have a radius of $0\aas@@fstack{\prime\prime}6$ for L1, L3 and the arc, and $0\aas@@fstack{\prime\prime}3$ for L2. Right panels: corresponding 1D spectra (in gray) in the observed frame, with absorptions lines indicated, as well as smoothed spectra (in color). The estimated redshifts of each component is also indicated; we measure it for L1 and L3, assume that L2 is at the same redshift as L1, and the source redsfhit has been measured in Kamieneski et al. (2023) based on radio observations. The hatched region indicates the gap caused to laser guiding. In this work we only use L1 and L3 spectra to extract stellar kinematics measurements, as L2 is highly blended with L1 and the signal from the arc is too faint.
3.1 Velocity dispersion measurement
@@
We measure the line-of-sight (LOS) central velocity dispersion from the MUSE observations of Caminha et al. (2023). We first extract integrated spectra for the three deflectors within circular apertures as shown on Fig. 2. The aperture radius is set to $0\aas@@fstack{\prime\prime}6$ which approximately corresponds to the MUSE resolution. For comparison, we also show one dimensional spectra corresponding to the brightest part of the arc, as well as a tentative extraction for L2. For the latter we use an aperture twice as small in order to reduce contamination by L1. We note that we only attempt to measure stellar kinematics of L1 and L3, since L2 would require a proper separation from L1, and the arc is too faint.
@@
We use the spectral template fitting software package pPXF (Cappellari & Emsellem 2004; Cappellari 2017) to model L1 and L3 spectra to measure their line-of-sight central velocity dispersion $\sigma_{v,\rm los}$ and redshift. As the signal-to-noise (S/N) of the spectra are significantly lower below 6000 $\AA$, we consider only wavelength interval $6650\,\AA-9300\,\AA$ for all kinematics measurements. We run pPXF multiple times in order to marginalize over possible systematic effects due to specific modeling assumptions. First, we vary the orders of the additive polynomial that captures large-scale variations in the continuum, from no polynomial up to order 8 polynomials. Second, we repeat the measurements (with all polynomial orders) varying this time the stellar template library. We use the Indo-US library (Valdes et al. 2004), which has a spectral resolution (FWHM) of $\sigma_{\rm temp,indo}=0.57$$\AA$ (for a dispersion of $0.4$ $\AA$), and the MILES library (Sánchez-Blázquez et al. 2006), which has a spectral resolution (FWHM) of $\sigma_{\rm temp,miles}=1.07$$\AA$ (for a dispersion of $0.9$ $\AA$). In empty patches in the vicinity of the lens, we fit Gaussian profiles to 7 sky emission lines to estimate the line spread function (LSF) and find $\sigma_{\rm lsf,obs}=1.34$$\AA$ (for a dispersion of $1.25$ $\AA$). Assuming a redshift of $z=0.87$, this leads to a rest-frame value of $\sigma_{\rm lsf,rest}=0.72$$\AA$, which we use as the spectral resolution of the MUSE observations. As the Indo-US library has a better resolution as our spectra, we need to convolve the spectral templates with a kernel of squared width $\sigma^{2}_{\rm diff}=\sigma^{2}_{\rm lsf,rest}-\sigma^{2}_{\rm temp,indo}$ and run the optimization to obtain the best-fit $\sigma_{v,\rm los}$ value. We apply a similar treatment for measurements obtained with the MILES library; however, these templates have a slightly lower resolution compared to our spectra, thus we apply the correction after the pPXF fit in order to retrieve the corrected best-fit $\sigma_{v,\rm los}$ value (see Eq. 5 of Cappellari 2017). Since pPXF requires an estimation of the noise per wavelength, and the variance from the MUSE data cube is unreliable (see e.g., Sect. 3.1.5 of Bacon et al. 2017), we follow pPXF examples [^2] https://github.com/micappe/ppxf_examples and assume a uniform error equal to unity throughout the entire wavelength range. We emphasize that this choice of noise level does not impact the best-fit values, but only error estimates based directly on the $\chi^{2}$, which is re-estimated after the fit. To perform all the aforementioned processing steps and measurements, we use an extended version of the pPXF wrapper code veldis (Mozumdar et al. 2023).
@@ 3.2 Contribution from the cluster
[…]
@@ 4.2 Point spread function
Figure 4: Point spread function (PSF) models for the main NIRCam filters used in this work, constructed from stars detected in the field. Each cutout has been truncated to 30 times the measured resolution (FWHM) in the corresponding filter. The angular size of the cutouts ranges from approximately $2^{\prime\prime}$ (F150W) to $5^{\prime\prime}$ (F356W).
@@ 4.3 Differentiable multi-band lens modeling
As the redshifts of L1 and L3 are slightly different (0.864 and 0.872), we check if these deflectors should be placed on different lens planes. We use the multi-lens plane interface of Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021) to estimate the error on the deflection field when assuming a single lens plane at the redshift of L1. More specifically, we compute the maximal deflection angle difference between single-plane and multi-plane deflection fields over the arc region. We find that the deflection angle changes only by $0\aas@@fstack{\prime\prime}0098$, which is $6.3$ times smaller than the JWST resolution in band F150W (${\rm FWHM_{\rm F150W}}=0\aas@@fstack{\prime\prime}062$ from Table 1 of Windhorst et al. 2023). Therefore, we do not employ the multi-lens plane formalism in this work, and essentially assume that L1, L2 and L3 are in the same lens plane.
@@
We model the surface brightness of the deflectors using a combination of Sérsic profiles (Sérsic 1963). The Sérsic profile is parametrized by its half-light radius $\theta_{\rm eff}$, Sérsic index $n_{\rm s}$, axis ratio $q_{\rm\ell}$, position angle $\phi_{\rm\ell}$, centroid $(x_{\rm\ell},y_{\rm\ell})$ and amplitude $I_{\rm eff}$ at $\theta_{\rm eff}$. We additionally include a uniform light profile to the lens light model in order to take into account potentially over- or under-subtracted sky level, parametrized by a single intensity $I_{\rm bkg}$.
@@ 4.4 Constraints from stellar kinematics
$\theta_{\rm E,SIS}$ [ ${}^{\prime\prime}$ ], realisticfootnotemark: $1.05\pm 0.05$ footnotemark: $0.49\pm 0.05$
$\theta_{\rm E,SIS}$ [ ${}^{\prime\prime}$ ], conservativefootnotemark: $0.86\pm 0.04$ footnotemark: $0.45\pm 0.05$
@@
\$$\star$\$\$\$$\star$\$\$footnotetext: Note that we infer that the density profile L1 is steeper than for an SIS from lensing constraints (Sect. 5.3).
@@
where $c$ is the speed of light, $D_{\rm s}$ and $D_{\rm s}$ are angular diameter distances to the source and between the deflector and the source, respectively. We add a multiplicative factor $f_{\rm cluster}$ to take into account the cluster contribution to the gravitational potential. As discussed in Sect. 3.2, the cluster realistically contributes to the mass of L1 and L3 at the $11\%$ and $9\%$ levels, respectively, hence we set $f_{\rm cluster,L1}=1-0.11$ and $f_{\rm cluster,L3}=1-0.09$. For angular diameter distances, we compute $D_{\rm s}$ and $D_{\rm ds}$ corresponding to L1 and L3 using redshift estimates from pPXF fits (i.e., $z_{\rm L1}=0.872$ and $z_{\rm L3}=0.864$), and assume $z_{\rm s}=2.291$ for the source redshift. The resulting probability distributions lead to $\theta_{\rm E,L1}=1\aas@@fstack{\prime\prime}05\pm 0\aas@@fstack{\prime\prime}05$ and $\theta_{\rm E,L3}=0\aas@@fstack{\prime\prime}49\pm 0\aas@@fstack{\prime\prime}05$, as shown in Fig. 5 (also compared to the distributions for a more conservative cluster contribution). We summarize in Table 1 the redshift, velocity dispersion and Einstein radius estimates. We emphasize that, to complement lensing constraints, we only use the Einstein radius estimate for L3 consistently with the assumption of an isothermal density profile used in the lens model. As we show in Sect. 4.3, we find that the mass density slope of L1 is steeper than isothermal.
@@ 4.5 Bayesian framework
Figure 6: Top row: example exposure maps for filters F200W and F [^4] W. The non-uniform exposure time is due to the combination of dithered exposures and the masking of cosmic rays. White isophotes indicate the position of the arcs and deflectors with respect to features in the exposure map. Bottom row: diagonal of the data covariance matrix (i.e., the noise map) estimated from the exposure map and the model-predicted flux to compute the shot noise contribution. The detector edge and cosmic ray imprints are still clearly visible in the noise map and are taken into account during modeling.
@@ 4.6 Modeling sequence
We start by modeling the surface brightness of the three deflectors L1, L2 and L3. Based on the NIRCam imaging data these are likely elliptical galaxies, although L2 and L3 have high ellipticities which could be the hint of a disk component. As all three deflectors are very bright and display significant deviations from axisymmetry, several model components a required to model their light distribution to acceptable levels. In particular we use 2 concentric Sérsic profiles for L1 and L2, and 3 concentric Sérsic profiles for L3. We also include a free constant background light component to properly account for any systematic offset from previous background subtraction steps. To maximize the fit quality in each of the six JWST filters and given the limited flexibility of the Sérsic profiles, we optimize lens light parameters in each bands separately. This single-band fitting step also allows us to measure the scatter in the position of the deflectors light distribution, and further use it as a prior for the centroids of their mass distribution. For each filter, we carefully design a mask to exclude the region containing significant source flux from the likelihood. We also exclude several other luminous objects in the field of view, mostly small galaxies or potential tidal stripping features. The optimization of all parameters is performed using the second-order gradient descent algorithm BFGS (Nocedal & Wright 2006), accessible in Herculens as a wrapper for the python package JaxOpt (Blondel et al. 2021). All lens light parameters are then fixed to their best-fit values for subsequent modeling steps.
@@
We then replace this intermediate source model with the 3D correlated field described in Eq. 1. The arc mask is updated such that it encloses all the flux from the source galaxy (importantly, it should also contain areas where its flux decreases to zero). Given a set of mass model parameters, this arc mask dynamically sets the extent of the pixelated grid in the source plane, such that the source grid pixel scale relative to the angular scale of the source remains approximately constant [^9] A similar strategy is used in the lens modeling software GLEE (Suyu & Halkola 2010; Suyu et al. 2012).. Beside the arc mask, the overall likelihood mask remains the same as for the lens light fitting step. We converge on a fiducial number of source pixels $n_{\rm src}=100$ along each spatial dimension, by progressively increasing $n_{\rm src}$ until we observe no significant changes in the residuals (i.e., the reduced $\chi^{2}$ remains approximately constant). We also find that increasing $n_{\rm src}$ does not significantly affect the morphology of the lensed galaxy, but only allows to capture more of the compact clumps within its disk, which only cover a negligible fraction of pixels entering the $\chi^{2}$ calculation. We maintain the lens light parameters fixed to their best-fit values, and jointly fit for the source and lens mass parameters. To allow our model to compensate for imperfect lens light modeling, we add a free uniform amplitude over the arc mask region, independently in each band.
@@
The lens models considered in this work have $\sim 10^{5}$ parameters [^10] The effective number of model parameters, in particular due to the field regularization of the source model, is much lower (although it is not explicit). and are constrained by $\sim 10^{6}$ pixels (excluding pixels outside the masked areas). We note that just-in-time compilation, gradient-based inference algorithms and GPU parallelization all lead to a low computation time. In particular, it takes approximately 140 minutes on a single NVIDIA A100 GPU to obtain one variational inference estimate of the joint posterior distribution, for a given model variation. Increasing the resolution of the source, even significantly, only marginally impacts runtime.

### v2

@@ 3 Stellar kinematics measurements
Figure 2: Extraction of MUSE aperture spectra within circular apertures, for L1, L2, L3 and the western part of El Anzuelo arc. Top left panel: luminosity-weighted median stack of the MUSE datacube, with contours from NIRCam/F [^4] W observations. Bottom left panel: MUSE median stack with circular apartures used to extract spectra. Circular apertures have a radius of $0\aas@@fstack{\prime\prime}6$ for L1, L3 and the arc, and $0\aas@@fstack{\prime\prime}3$ for L2. Right panels: corresponding 1D spectra (in gray) in the observed frame, with absorptions lines indicated, as well as smoothed spectra (in color). The hatched region indicates the gap caused to laser guiding. The estimated redshift of each component is also indicated in the upper left corner. We measure the redshifts for L1 and L3, but assume that L2 is at the same redshift as L1 and that the source is at redshift $z=2.291$ (Kamieneski et al. 2023, using a CO line detection in ALMA observations). Some key absorption lines are indicated by thin dashed gray lines (line labels are indicated in top and bottom panels only to avoid clutter). For the arc, we indicate two sets of lines corresponding to our assumed redshift (black and dashed lines) and the alternative possible redshift $z=3.388$ (gray and dotted lines, for details see Kamieneski et al. 2023). In this work we only use L1 and L3 spectra to extract their stellar kinematics measurements, as L2 is highly blended with L1 and the signal from the arc is too faint.
3.1 Redshift and velocity dispersion measurements
@@
We measure redshifts and line-of-sight (LOS) central velocity dispersions from the MUSE observations of Caminha et al. (2023). We first extract integrated spectra for the three deflectors within circular apertures as shown on Fig. 2. The aperture radius is set to $0\aas@@fstack{\prime\prime}6$ which approximately corresponds to the MUSE resolution. For comparison, we also show one dimensional spectra corresponding to the brightest part of the arc, as well as a tentative extraction for L2. For the latter we use an aperture twice as small in order to reduce contamination by L1. We note that we only attempt to measure stellar kinematics of L1 and L3 for which we can detect clear absorption features. The spectrum of L2 suggests that this galaxy lies at a redshift very close to L1. However, extracting robust stellar kinematics properties for L2 would require to properly separate the contributions from L1 and L2, which is outside the scope of this work.
Due to low signal-to-noise (S/N), we are unable to detect reliable features in the spectrum of the lensed arc, as shown in the bottom right panel of Fig. 2, such that we cannot robustly measure its redshift by fitting stellar templates. However, we indicate in Fig. 2 some absorption lines assuming the most probably redshift $z=2.291$ inferred by Kamieneski et al. (2023) assuming the CO line detection in ALMA data corresponds to the CO(3-2) transition. If instead it corresponds to the CO(4-3) transition, Kamieneski et al. (2023) mention that the redshift would change to $z=3.388$. For comparison, we indicate the spectral lines corresponding to this alternate redshift in Fig. 2 with dotted gray lines (bottom right panel). While the low S/N of the MUSE spectrum does not provide a definitive answer, the indicated absorption lines seem to align better with possible spectral features at $z=2.291$ compared to $z=3.388$. We note that the difference in angular-to-physical size between these two redshift estimates is only $0.8\,{\rm kpc}\,{\rm arcsec}^{-1}$. In the remaining of the work, we follow Kamieneski et al. (2023) and assume $z_{\rm s}=2.291$ for the source redshift, and we show in Appendix B that the alternative value for $z_{\rm s}$ impacts only marginally our results.
@@
We use the spectral template fitting software package pPXF (Cappellari & Emsellem 2004; Cappellari 2017) to model L1 and L3 spectra to measure their line-of-sight central velocity dispersion $\sigma_{v,\rm los}$ and redshift. As the S/N of the spectra are significantly lower below 6000 $\AA$, we consider only wavelength interval $6650\,\AA-9300\,\AA$ for all kinematics measurements. We run pPXF multiple times in order to marginalize over possible systematic effects due to specific modeling assumptions. First, we vary the orders of the additive polynomial that captures large-scale variations in the continuum, from no polynomial up to order 8 polynomials. Second, we repeat the measurements (with all polynomial orders) varying this time the stellar template library. We use the Indo-US library (Valdes et al. 2004), which has a spectral resolution (FWHM) of $\sigma_{\rm temp,indo}=0.57$$\AA$ (for a dispersion of $0.4$ $\AA$), and the MILES library (Sánchez-Blázquez et al. 2006), which has a spectral resolution (FWHM) of $\sigma_{\rm temp,miles}=1.07$$\AA$ (for a dispersion of $0.9$ $\AA$). In empty patches in the vicinity of the lens, we fit Gaussian profiles to 7 sky emission lines to estimate the line spread function (LSF) and find $\sigma_{\rm lsf,obs}=1.34$$\AA$ (for a dispersion of $1.25$ $\AA$). Assuming a redshift of $z=0.87$, this leads to a rest-frame value of $\sigma_{\rm lsf,rest}=0.72$$\AA$, which we use as the spectral resolution of the MUSE observations. As the Indo-US library has a better resolution as our spectra, we need to convolve the spectral templates with a kernel of squared width $\sigma^{2}_{\rm diff}=\sigma^{2}_{\rm lsf,rest}-\sigma^{2}_{\rm temp,indo}$ and run the optimization to obtain the best-fit $\sigma_{v,\rm los}$ value. We apply a similar treatment for measurements obtained with the MILES library; however, these templates have a slightly lower resolution compared to our spectra, thus we apply the correction after the pPXF fit in order to retrieve the corrected best-fit $\sigma_{v,\rm los}$ value (see Eq. 5 of Cappellari 2017). Since pPXF requires an estimation of the noise per wavelength, and the variance from the MUSE data cube is unreliable (see e.g., Sect. 3.1.5 of Bacon et al. 2017), we follow pPXF examples [^2] https://github.com/micappe/ppxf_examples and assume a uniform error equal to unity throughout the entire wavelength range. We emphasize that this choice of noise level does not impact the best-fit values, but only error estimates based directly on the $\chi^{2}$, which is re-estimated after the fit. To perform all the aforementioned processing steps and measurements, we use an extended version of the pPXF wrapper code veldis (Mozumdar et al. 2023).
@@ 3.2 Contribution from the cluster
Since the cluster mass distribution might have a non-negligible effect on the kinematics of L1 and L3, we estimate its impact on $\sigma_{v,\rm los}$ for these two galaxies. Based on the model of Caminha et al. (2023), we compute the one dimensional projected total mass map without L1, L2 and L3 and measure the cluster mass contribution within the galaxies. Due to the lack of information about the cluster mass profile along the line of sight, we consider two approximations. The first consists of assuming that all cluster mass and galaxies are located at a thin plane. In this case, we use a $0\aas@@fstack{\prime\prime}6$ aperture centered on L1 and L3 (consistently with the measurement aperture) to compute the cluster mass. This estimate is conservative and potentially overestimates the cluster contribution. The second method assumes a mass profile along the line of sight given by a cored isothermal profile (see e.g., Elíasdóttir et al. 2007), that is $\rho(z)\propto 1/(r_{\rm core}^{2}+z^{2})$, where $r_{\rm core}$ is the core radius of the main cluster halo. In this case, the cluster mass is integrated within the same aperture but out to two times the cut radii of L1 and L3 along the $z$ direction. Such assumption of a symmetric profile along the line of sight is still an approximation, but gives a more realistic estimation of the impact on the galaxies kinematics. We obtain that the additional mass of the cluster component increases the galaxies velocity dispersion by factors of 27% (conservative) or 11% (realistic) for L1, and only 16% or 9% for L3.
@@ 4.2 Point spread function
Figure 4: Point spread function (PSF) models for the main NIRCam filters used in this work, constructed from stars detected in the field. Each cutout has been truncated to 30 times the measured resolution (FWHM) in the corresponding filter. The angular size is indiciated by a scale bar at the bottom right of each panel.
@@ 4.3 Differentiable multi-band lens modeling
As the redshifts of L1 and L3 are slightly different (0.872 and 0.864, respectively), we check if these deflectors should be placed on different lens planes. We use the multi-lens plane interface of Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021) to estimate the error on the deflection field when assuming a single lens plane at the redshift of L1. More specifically, we compute the maximal deflection angle difference between single-plane and multi-plane deflection fields over the arc region. We find that the deflection angle changes only by $0\aas@@fstack{\prime\prime}0098$, which is $6.3$ times smaller than the JWST resolution in band F150W (${\rm FWHM_{\rm F150W}}=0\aas@@fstack{\prime\prime}062$ from Table 1 of Windhorst et al. 2023). Therefore, we do not employ the multi-lens plane formalism in this work, and essentially assume that L1, L2 and L3 are in the same lens plane.
@@
We model the surface brightness of the deflectors using a combination of Sérsic profiles (Sérsic 1963). The Sérsic profile is parametrized by its half-light radius $\theta_{\rm eff}$, Sérsic index $n_{\rm s}$, axis ratio $q_{\rm\ell}$, position angle $\phi_{\rm\ell}$, centroid $(x_{\rm\ell},y_{\rm\ell})$ and amplitude $I_{\rm eff}$ at $\theta_{\rm eff}$. We additionally include a uniform light profile to the lens light model in order to take into account potentially over- or under-subtracted sky level, parametrized by independent intensities $I_{\rm bkg}$ in each band.
@@ 4.4 Constraints from stellar kinematics
$\theta_{\rm E,SIS}$ [ ′′ ], realisticfootnotemark: $1.05\pm 0.05$ footnotemark: $0.49\pm 0.05$
$\theta_{\rm E,SIS}$ [ ′′ ], conservativefootnotemark: $0.86\pm 0.04$ footnotemark: $0.45\pm 0.05$
@@
\$$\star$\$\$\$$\star$\$\$footnotetext: In this work, we infer that the density profile L1 is steeper than for an SIS from lensing constraints (Sect. 5.3).
@@
where $c$ is the speed of light, $D_{\rm s}$ and $D_{\rm ds}$ are angular diameter distances to the source and between the deflector and the source, respectively. We add a multiplicative factor $f_{\rm cluster}$ to take into account the cluster contribution to the gravitational potential. As discussed in Sect. 3.2, the cluster realistically contributes to the mass of L1 and L3 at the $11\%$ and $9\%$ levels, respectively, hence we set $f_{\rm cluster,L1}=1-0.11$ and $f_{\rm cluster,L3}=1-0.09$. For angular diameter distances, we compute $D_{\rm s}$ and $D_{\rm ds}$ corresponding to L1 and L3 using redshift estimates from pPXF fits (i.e., $z_{\rm L1}=0.872$ and $z_{\rm L3}=0.864$), and assume $z_{\rm s}=2.291$ for the source redshift. The resulting probability distributions lead to $\theta_{\rm E,L1}=1\aas@@fstack{\prime\prime}05\pm 0\aas@@fstack{\prime\prime}05$ and $\theta_{\rm E,L3}=0\aas@@fstack{\prime\prime}49\pm 0\aas@@fstack{\prime\prime}05$, as shown in Fig. 5 and compared to the distributions for a more conservative cluster contribution. We summarize in Table 1 the redshift, velocity dispersion and Einstein radius estimates.
We emphasize that, to complement lensing constraints, we only use the Einstein radius estimate for L3 consistently with the assumption of an isothermal density profile used in the lens model. As shown in Kamieneski et al. (2023) and in Sect. 4.3, the lensing data is not constraining enough to probe the mass distribution of L3 beyond the assumption of an isothermal profile. For L1 however, we find 4.3 that the mass density slope at the Einstein radius is steeper than isothermal.
@@ 4.5 Bayesian framework
Figure 6: Non-uniform exposure time over the field of view of El Anzuelo due to the combination of dithered exposures and masking of cosmic rays. Top row: example exposure maps for filters F200W and F [^4] W. White isophotes indicate the position of the arcs and deflectors with respect to features in the exposure map. Bottom row: diagonal of the data covariance matrix (i.e., the noise map) estimated from the exposure map and the model-predicted flux to compute the shot noise contribution. The detector edge and cosmic ray imprints are still clearly visible in the noise map and are taken into account during modeling.
@@ 4.6 Modeling sequence
We start by modeling the surface brightness of the three deflectors L1, L2 and L3. Based on the NIRCam imaging data these are likely elliptical galaxies, although L2 and L3 have high ellipticities which could be the hint of a disk component. As all three deflectors are very bright and display significant deviations from axisymmetry, several model components are required to model their light distribution to acceptable levels. In particular we use 2 concentric Sérsic profiles for L1 and L2, and 3 concentric Sérsic profiles for L3. We also include a free constant background light component to properly account for any systematic offset from previous background subtraction steps. We do not model possible intra-cluster light (ICL), which is clearly visible only closer to the main cluster components in the JWST data. El Anzuelo is located at a projected distance of $\sim 230$ kpc from the brightest cluster member (BCG), in a region where the ICL is fainter. Moreover, as estimated by Diego et al. (2023), at a projected distance of 100 kpc from the BCG, the ICL in El Gordo contributes to $\lesssim 1\%$ of the projected mass. If a significant fraction of ICL was missing in our model, it would appear in the model residuals as a large scale gradient over the field of view, consistently between the filters. As we do not observe such large scale residuals (see also Appendix C.1), we do not further increase the flexibility of our light model. To maximize the fit quality in each of the six JWST filters and given the limited flexibility of the Sérsic profiles, we optimize lens light parameters in each bands separately. This single-band fitting step also allows us to measure the scatter in the position of the deflectors light distribution, and further use it as a prior for the centroids of their mass distribution. For each filter, we carefully design a mask to exclude the region containing significant source flux from the likelihood. We also exclude several other luminous objects in the field of view, mostly small galaxies or potential tidal stripping features. The optimization of all lens light parameters is performed using the second-order gradient descent algorithm BFGS (Nocedal & Wright 2006), accessible in Herculens as a wrapper for the python package JaxOpt (Blondel et al. 2021). All lens light parameters are then fixed to their best-fit values for subsequent modeling steps.
@@
We then replace this intermediate source model with the 3D correlated field described in Eq. 1. The arc mask is updated such that it encloses all the flux from the source galaxy, such that it covers areas that contain no source flux as well (ensuring that the full extent of the source is being reconstructed). Given a set of mass model parameters, this arc mask dynamically sets the extent of the pixelated grid in the source plane, such that the source grid pixel scale relative to the angular scale of the source remains approximately constant [^9] A similar strategy is used in the lens modeling software GLEE (Suyu & Halkola 2010; Suyu et al. 2012).. Beside the arc mask, the overall likelihood mask remains the same as for the lens light fitting step. We converge on a fiducial number of source pixels $n_{\rm src}=100$ along each spatial dimension, by progressively increasing $n_{\rm src}$ until we observe no significant changes in the residuals (i.e., the reduced $\chi^{2}$ remains approximately constant). We also find that increasing $n_{\rm src}$ does not significantly affect the morphology of the lensed galaxy, but only allows to capture more of the compact clumps within its disk, which only cover a negligible fraction of pixels entering the $\chi^{2}$ calculation. We maintain the lens light parameters fixed to their best-fit values, and jointly fit for the source and lens mass parameters. To allow our model to compensate for imperfect lens light modeling, we add a free uniform amplitude over the arc mask region, independently in each band.
@@
The lens models considered in this work have $\sim 10^{5}$ parameters [^10] The “effective” number of model parameters, in particular due to the field regularization of the source model, is much lower (although it is not explicit). and are constrained by $\sim 10^{6}$ pixels (excluding pixels outside the masked areas). We note that just-in-time compilation, gradient-based inference algorithms and GPU parallelization all lead to a low computation time. In particular, it takes approximately 140 minutes on a single NVIDIA A100 GPU to obtain one variational inference estimate of the joint posterior distribution, for a given model variation. Increasing the resolution of the source, even significantly, only marginally impacts runtime.